• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name LeRose

    Log entry time 19:34:15 on July 01, 2010

    Entry number 329320

    keyword=Partial Shift Summary

    Day shift 7/1/2010
    John LeRose (SL)
    Al Odian (TO)
    
    At shift start we are in restricted access.
    We've been pumping on the target chamber since 14:30
    Setting up right arm for 1.131 GeV (Q2/3 cycled to 1000 Amps)
       Had problems with the right arm NMR(seehalog 329316 for solution).
       In the end successfully setup for 1.131 GeV/c (cycled quads)
    
    19:30 The plan for the evening is to continue DAQ work in the Hall
    	while waiting for the target vacuum to drop.
    
    I'm going home, call me if you need me (810 6784)